# Fairly Accurate: Fairness-aware Multi-group Target Detection in Online Discussion

### v2

@@ 10 Conclusion
We have proposed a fairness-aware formulation of multi target-group detection, focusing on realistic scenarios where social media posts may target multiple demographic groups. Motivated by the symmetric harm of misclassification in this setting, we adopted Accuracy parity (AP) as a fairness criteria and proposed $GAP_{multi}$, a loss function that minimizes bias in a multi-group setting. Our method jointly optimizes for group fairness and utility, aligning training dynamics with evaluation measures. Through theoretical analysis, we demonstrated the incompatibility between Accuracy Parity and Equalized Odds, revealing trade-offs that are important to consider in group identity sensitive NLP tasks. Empirical evaluations on the MHS corpus, a large multi-label dataset with fine-grained demographic annotations, show that our approach consistently reduces bias across groups while retaining overall utility. By providing the first evaluation for this task, our work advances model development in fairness-aware target-group detection, contributing towards the development of safer and inclusive online spaces.
@@ Appendix A Impossibility Proofs
Assume two groups $A$ and $B$. Let the number of positive and negative examples in group $A$ be $P_{A}$ and $N_{A}$ respectively. Similarly, $P_{B}$ and $N_{B}$ for group $B$ respectively. Proportion of positive to negative examples in a group, i.e., $\delta_{A}=\frac{P_{A}}{N_{A}}$ and $\delta_{B}=\frac{P_{B}}{N_{B}}$.
@@
For Group A, we have the following:
@@
Special Case 2: Base rates are not equal, but $TPR$ and $FPR$ sum to one. $TPR+FPR=1$. Under this special setting of unequalized base rates, Eq. 12 and Eq. 13 becomes the same. This is equivalent to random predictions by the classification model.
@@
Proof of Theorem 2. The False Positive and Negative Error Disparity (FPNED) [7] can be broken down into FPED and FNED.
@@ Appendix B Experimental Details
B.1 Setup
Experiments use a Nvidia 2060 RTX Super 8GB GPU, Intel Core i7-9700F 3.0GHz 8-core CPU and 16GB DDR4 memory. We use the Keras [8] library on a Tensorflow 2.0 backend with Python 3.7 to train the networks in this paper. For optimization, we use AdaMax [25] with parameters (lr=0.001) and $1000$ steps per epoch. We also setup an EarlyStopping criteria in case the with a min_delta of 1e-4 and patience of 5.
B.2 Data Pre-processing
From the set of all posts, we filter out ones that target at least one or more demographic group(s), resulting in a dataset size of 44,816 posts. From there we perform a 80%-20% split for training vs. testing data, for all cluster sizes that allowed for a SkLearn’s [36] stratified sampling following the label split ratio. This resulted in a train set of size 36,412 posts and test set of size 8,404 posts.
B.3 Loss
@@
Appendix C Pairwise Group Classifiers
@@
GAP 92.34 91.79 92.06 0.55
Table 7: Optimizing across groups results in improving overall error, indicating that the group label provides an extra dimension for the loss to stabilize at a better local optima.
@@
GAP 89.65 90.82 90.23 1.17
Table 8: Performance varies across groups due to their population size, where the statistically major group dominates.
@@
GAP 88.36 86.23 87.29 2.13
Table 9: Groups have similar population, but performance varies due to one group being more difficult to model.

### v3

@@ 10 Conclusion
We have motivated and proposed a framework for fairness-aware multi-group target detection, focusing on realistic scenarios where social media posts may target multiple demographic groups simultaneously. Given the symmetric harm of detection errors in this setting, we adopted Accuracy Parity (AP) as a fairness criteria and proposed the $GAP_{multi}$ loss function to minimize bias in a multi-group setting. We jointly optimize for both group fairness and utility, aligning training dynamics with evaluation measures in a robust and scalable manner. We also present theoretical analysis showing the incompatibility between AP and Equalized Odds (EO). Our empirical results show consistently reduced bias across groups while retaining overall utility on fairness-aware multi-group target detection contributes towards developing safer and more inclusive online spaces.
11 Ethical Considerations Statement
In safety-critical domains, such as toxicity detection and content moderation, the ability to identify specific groups serves a vital protective function by shielding vulnerable populations from harm and delivering a clear societal value that justifies its use. Conversely, in contexts like targeted advertising, the same technical capability could enable profiling or manipulation without stringent oversight, posing significant ethical risks. Researchers should acknowledge potential risks alongside potential benefits of work in this area, and to make every effort to ensure that potential benefits outweigh potential risks. To this end, we explicitly recommend that practical deployments prioritize protective applications, incorporate transparent ethical guidelines, and implement rigorous impact assessments. Moreover, users should always be informed of when and why target-group detection systems are deployed. This approach strives to ensure and support our goal that the tool aligns with societal values and mitigates the risk of exacerbating existing inequalities.
Our experiments used two existing, public datasets, the MHS corpus [35] and HateXplain [29]. For MHS, its authors note that their data annotation process was approved by UC Berkeley’s IRB. Their annotators were allowed to omit any demographic information, and all data samples were also anonymized to protect annotator privacy. For HateXplain, its authors note that posts were anonymized by replacing the usernames with <user> token. Our own experiments use only the target’s demographic-group flag and we do not infer any other user/annotator sensitive information.
@@ Appendix A Impossibility Proofs
Assume two groups $A$ and $B$. Let the number of positive and negative examples in group $A$ be $P_{A}$ and $N_{A}$ respectively. Similarly, $P_{B}$ and $N_{B}$ for group $B$. Proportion of positive to negative examples in a group, i.e., $\delta_{A}=\frac{P_{A}}{N_{A}}$ and $\delta_{B}=\frac{P_{B}}{N_{B}}$.
@@
For Group A and Group B, we have the following:
@@
Special Case 2: Base rates are not equal, but $TPR$ and $FPR$ sum to one. $TPR+FPR=1$. Under this special setting of unequalized base rates, Eq. 7 and Eq. 8 becomes the same. This is equivalent to random predictions by the classifier.
@@
Proof of Theorem 2. The False Positive and Negative Equality Difference (FPNED) [6] can be broken down into FPED and FNED.
@@ Appendix B Experimental Details
Setup
Experiments use a Nvidia 2060 RTX Super 8GB GPU, Intel Core i7-9700F 3.0GHz 8-core CPU and 16GB DDR4 memory. We use Keras [7] library on a Tensorflow 2.0 backend with Python 3.12 to train the networks in this paper. For optimization, we use AdaMax [23] with parameters (lr=0.001) and $1000$ steps per epoch. We also setup an EarlyStopping criteria in case the with a min_delta of $1\textsc{e-}4$ and patience of 5. The random seed was set to 42 to avoid varaibility.
Optimizer
We choose AdaMax (infinity-norm variant of Adam) due to its faster convergence from fresh starts, i.e., model training from random weights. It is also more stable than Adam in our case with sparse or unbalanced gradients. This makes tuning the learning rate less fragile, which is convenient when balancing $\lambda$ for fairness.
Regularization
[…]
Dataset
We use the MHS corpus [35] via the HuggingFace library [13]. Its 135,556 posts and seven race groups: Asian, Black, Latinx, Middle-Eastern, Native-American, Pacific-Islander and White. Fig. 5(a) shows how frequently each group is targeted. Fig. 5(b) shows the number of groups targeted per post [^2] For groups of size six and seven, posts target a broader range of demographic categories, often using inclusive terms like ‘non-white’, as shown in Table 1. This explains the higher frequency of such posts compared to those targeting fewer than five groups..
(a) Number of posts targeting each group.
(b) Number of posts targeting multiple groups.
We also use the HateXplain [29] dataset containing 57,687 posts sourced from Twitter and Gab with the following demographic groups: African, Arab, Asian, Caucasian, and Hispanic, with African as the statistical majority group ($\sim$9k posts), while Asian and Hispanic are statistical minorities. Fig. 6(a) shows how frequently each group is targeted. Fig. 6(b) shows the number of groups targeted per post.
(a) Number of posts targeting each group.
(b) Number of posts targeting multiple groups.
Figure 6: Summary statistics of the HateXplain [29] show the distribution of posts targeting demographic groups. The African community is the statistical majority, while Native American and Pacific Islander are statistical minorities. Additionally, the dataset includes posts targeting multiple groups, reflecting its multi-group nature.
Dataset Pre-processing
From the set of all posts across both datasets, we filter ones that target at least one or more demographic group(s). In MHS Corpus this results in a dataset size of 44,816 posts. We perform a 70-10-20% stratified split for training, validation and testing data. HateXplain’s filtered posts has 15,410 samples, which comes with its own pre-defined 80-10-10% training, validation and testing split. We further divide the data into clusters by number of targeted groups, for all cluster sizes that allowed for a SkLearn’s [34] stratified sampling on the label split ratio.
Loss Function
@@
Appendix C Additional Experiments
C.1 Evaluation Measures
Balanced Accuracy (BA)
Unlike standard accuracy, which can be misleading in the presence of label imbalance, Balanced Accuracy measures accuracy by averaging per-label performance, addressing dataset imbalance. By considering both the (TPR: true positive rate) and (TNR: true negative rate) of each label, BA effectively captures the model’s ability to correctly classify instances across all labels, regardless of their prevalence.
$BA=(TPR+TNR)/2.0$ (10)
Avg. BA
We report the macro averaged BAs as a summary statistic, to equally weigh performance across all groups. The Avg. BA (macro) treats each group equally, ensuring that the classifier’s performance is evaluated in a balanced manner across all demographic groups. macro BA takes precedence over micro BA due to its emphasis on equal representation of groups and and fairness. Unlike micro BA, which heavily weights larger groups, macro BA
$Avg.\,BA=\frac{1}{G}\sum_{g=1}^{G}BA(g)$ (11)
Hamming Loss (HL)
This is a widely used metric for assessing the performance of multi-label classifiers. For a dataset with $N$ instances and $G$ labels, the Hamming Loss (HL) measures the average fraction of incorrectly predicted labels across all instances, with $hamming(y_{i}(g),\hat{y}_{i}(g))$ as an indicator function of 1 if the $g$-th label for instance $i$ is incorrectly predicted and 0 otherwise. Specifically, it measures the average fraction of labels that are misclassified in comparison to the true label set. HL and Subset Accuracy Loss are comparable under small label cases [41], hence we just report HL.
C.2 Results on MHS
To assess loss function performance consistency across different text encoder representations, we also evaluate with RoBERTa-base [27] instead of DistilBERT. RoBERTa-base did not yield any significant gain in utility, besides adding to runtime due to its larger size. Results shown in Table 8 validate that even with upgraded text encoder model, $GAP_{multi}$ achieves the minimum group disparity (Max. Diff.) performance among the compared losses.
Table 8: Evaluation of loss function on the MHS corpus using RoBERTa-base as the text encoder.
MHS Corpus - Balanced Accuracy (BA) ($\uparrow$)
[…]
For $\lambda$-tuning on MHS, see Fig. 7 for sensitivity of accuracy-fairness (Average Balanced Accuracy - Maximum Difference) vs. $\lambda$. The stopping criteria for this experiment was triggered at $\lambda$=0.4, validation accuracy stabilized at 81.84%. Note that the trajectory of the curve goes beyond traditional optimization (Eq. 3) where with increasing fairness (lower bias), overall accuracy should be correspondingly decrease. However, as we discussed for MHS, both CLA and $GAP_{multi}$ have improvements in accuracy possibly due to the extra group dimension that is invoked during optimization. Hence for this specific data, we have empirically gained in both accuracy and group fairness with increasing $\lambda$.
Figure 7: Hyperparameter results showing variation of Avg. BA and Max. Diff. vs. $\lambda$ on the MHS corpus. The extra group dimension provides both CLA and $GAP_{multi}$ an extra dimension to optimize over, hence we empirically gain improvements in both vertical axes.
C.3 Results on HateXplain
Fig. 8 shows the visualization corresponding to Table 4 in the main material on the 5-group HateXplain. For this dataset, CLA achieves the highest Avg. BA (74.84), but both OE and $GAP_{multi}$ achieve statistically comparable performance. Meanwhile, $GAP_{multi}$ achieves a statistically significant reduction in Max. Diff.. The comparative reduction in disparity across groups, alongside similar Avg. BA, shows the method’s effectiveness as a fair and utility-preserving loss function for multi target-group identification task.
Figure 8: Visualization of the BA values achieved by each loss over the 5 demographic groups in HateXplain, using DistilBERT as the text encoder. The maximum difference (Max. Diff.) of BA achieved for each loss across groups is also shown. $GAP_{multi}$ performs best with lowest Max Diff. of 5.2, showcasing that it greatly reduces the disparities in performance across groups vs. other losses.
Figure 9 displays pairwise absolute differences of Balanced Accuracy (BA) across various groups ($|\textit{left - bottom}|$) in HateXplain. Higher heatmap values indicate classifier bias, revealing consistent inequities, particularly for overall error (OE) rates, where the (African, Asian) pair has the largest Max. Diff. (11.0). While CLA (optimized for FNR) improves over OE, its misaligned objective (minimizing false negatives, rather than equitable performance), results in a heatmap in-between that of OE and $GAP_{multi}$. In contrast, the $GAP_{multi}$ loss, explicitly designed for balanced group performance, shows substantially fewer extremes (Max. Diff.: 5.2), with smoother heatmap transitions indicating equitable classifier outcomes, aligning with our fair target-group detection objective.
Figure 9: Heatmap of pairwise absolute difference of BA across groups in test set (HateXplain) as an indicator for bias and disparate impact. OE has the highest performance gap (Max Diff = ) across groups as indicated by the extremes of color, not only across one group- pair but consistently across multiple group pairs. $GAP_{multi}$ has the least spread in pairwise error values (Max. Diff. = 5.2), evident from the flatness of color, indicating least disparate impact across groups.
Fig. 10 shows sensitivity of accuracy-fairness (Average Balanced Accuracy - Maximum Difference) vs. $\lambda$. The stopping criteria for this experiment was triggered at $\lambda$=0.4, the bias criteria stabilized at 5.25, with accuracy 74.23%. Beyond that point accuracy keeps on slightly decreasing while the fairness criteria plateaus. Hence we stop at $\lambda$=0.4 since all the accuracy values to the right are dominated values, hence ignored. In contrast to what we observe for the MHS Corpus, this curve follows the standard optimization trend where decreasing bias comes at the expense of overall accuracy.
Figure 10: Hyperparameter Tuning Results showing variation of Avg. BA and Max. Diff. vs. $\lambda$ on HateXplain. For this dataset, with increasing values of $\lambda$ we observe a steady decrease in accuracy at the cost of bias.
Appendix D Pairwise Group Classifiers on MHS
@@
Table 9: Different 2-group setting highlighting the performance variation across the measures for different groups. (a) Optimizing across groups results in improving overall error, indicating that the group label provides an extra dimension for the loss to stabilize at a better local optima. (b) Performance varies across groups due to their population size, where the statistically major group dominates. (c) Groups have similar population, but performance varies due to one group being more difficult to model.
@@
$GAP_{multi}$ 92.34 91.79 92.06 0.55
(a) Case I
@@
$GAP_{multi}$ 89.65 90.82 90.23 1.17
(b) Case II
@@
$GAP_{multi}$ 88.36 86.23 87.29 2.13
(c) Case III